If 10 ampere current is flowing through primary coil and their transformation ratio is 20.Find the value of current flowing in secondary coil?
1
Expert's answer
2020-09-22T15:34:49-0400

Solution

Current in primary coil(Ip) =10A

Transformation ratio (K) =20

Relationship between currents and transformation ratio is given by

"\frac{I_p}{I_s}=K"

Therefore

Current in secondary coil

"I_s=\frac{I_p}{K}=\frac{10}{20}=0.5A"

So current in secondary coil is 0.5A.